How many minutes are needed by a car whose speed is 40.0 km/h to cover a distance of 800 m?
Answer by stanbon(75887)

How many minutes are needed by a car whose speed is 40.0 km/h to cover a distance of 800 m?
-----------------------------------------------------------------
40 km/hr = 40 (1000 meters/60 minutes)
= (2000/3) meters/minute
===============================
distance/rate = time
800 meters / (2000/3 meters/min)
---
= 2400/2000 min
= 1.2 minutes
=====================
Cheers,
Stan H.